The angular velocity of a rotating rigid body increases from 500 to 1500 rev/min in 120sec.
A, what is the angular acceleration of the body?
B, through what angle does it turn in this 120sec?
1
Expert's answer
2020-05-07T09:59:47-0400

a)


"\alpha=2\pi\frac{1500-500}{60\cdot 120}=0.873\frac{rad}{s^2}"

b)


"\theta=0.5\alpha t^2=0.5(0.873)(120)^2=6283\ rad"